# Specific heat study

### v1

@@ IV Discussion and summary
In this study, we presented the specific heat probes of several quantum three-dimensional Heisenberg magnets. Our main findings are reported in Figs. 2, 3, and 4. While ferromagnets and bipartite-lattice antiferromagnets can be straightforwardly investigated by applying quantum Monte Carlo simulations, frustrated-lattice antiferromagnets require more sophisticate techniques: We use high-temperature expansion series complemented by the entropy-method interpolation. Our findings illustrate effects of geometry on the ordering temperature (ferromagnets and bipartite-lattice antiferromagnets). For frustrated-lattice antiferromagnets, the obtained specific heat, which interpolates between assumed low-energy properties and rigorously known high-temperature ones, provides some information about a possible energy spectrum of the quantum spin system at hand. Although our study is pure theoretical, it may be useful from the solid-state perspective, if the corresponding compounds are available.
@@ Acknowledgements
Appendix A: Pyrochlore-lattice antiferroamgnet
Figure 6: Padé approximants to high-temperature expansion series for the $S=1/2$ Heisenberg antiferromagnet on pyrochlore lattice.
For the pyrochlore-lattice antiferromagnet the coefficients $d_{2},\ldots,d_{17}$ are reported in Ref. [14] (the first thirteen coefficients can be also found in Ref. [13]). Simple Padé approximants to high-temperature expansion series (6) are shown in Fig. 6. As can be seen from this figure, they can provide reliable results above $T\approx 0.7$.
To apply the entropy-method interpolation, we use the ground-state energy $e_{0}$ found in Refs. [21, 22] and make an assumption about the decay of the specific heat as $T\to 0$. We try a power-law decay with several exponents $\alpha=1,2,3$, as well as an exponential decay. However, none of the made assumptions yields an excellent temperature profile $c(T)$ with respect to large number $n_{\rm cP}$ of almost coinciding $c(T)$ which follow from $n_{\rm P}=u+d-3=14$ different $[u,d](e)$ approximant. On the other hand, we may try to determine $e_{0}$ as in Ref. [19]. However, we get $n_{\rm cP}=7$ for $e_{0}=-0.423\,4,\ldots,-0.420\,2$ if $\alpha=1$, $n_{\rm cP}=6$ for $e_{0}=-0.461\,3,\ldots,-0.42$ if $\alpha=2$, or $n_{\rm cP}=7$ for $e_{0}=-0.465\,2,\ldots,-0.457\,8$ if $\alpha=3$; these values are quite far from $e_{0}=-0.49$ [21] and $e_{0}=-0.477$ [22]. Apparently the calculation of $c(T)$ for the pyrochlore-lattice $S=1/2$ Heisenberg antiferromagnet is far from being finally settled.
Appendix B: Hyperkagome-lattice antiferromagnet
Figure 7: Entropy-method calculations for the $S=1/2$ Heisenberg antiferromagnet on hyperkagome lattice. Top: Padé approximants of high-temperature expansion series. Bottom: Ground-state energy determination within the entropy method.
We begin with the coefficients $d_{2},\ldots,d_{16}$ in Eq. (6), which are as follows:
$d_{2}{=}\frac{3}{8},\;d_{3}{=}0,\;d_{4}{=}-\frac{51}{128},\;d_{5}{=}0,\;d_{6}{=}\frac{349}{1\,024},$
$d_{7}{=}-\frac{7}{1\,280},\;d_{8}{=}-\frac{131\,449}{491\,520},$
$d_{9}{=}\frac{349}{35\,840},\;d_{10}{=}\frac{11\,103\,797}{55\,050\,240},$
@@
These coefficients can be found, in principle, in Ref. [15], where they are presented, however, as decimal fractions. In Eq. (Appendix B: Hyperkagome-lattice antiferromagnet) they are given as common fractions that maybe sometime more convenient.
The entropy method calculations are described in detail in Ref. [20], but now we use as an input 20 coefficients in Eq. (6) (instead of 16 in Ref. [20]). From the upper panel of Fig. 7 one concludes, that Padé approximants of high-temperature expansion series can give $c(T)$ for $T$ exceeding $\approx 0.3$. In the lower panel of Fig. 7 we closely follow Ref. [20] and present the number of almost coinciding $c(T)$ denoted as $n_{\rm cP}$ among all considered $n_{\rm P}=u+d-3$ Padé approximants $[u,d](e)$ as a function of the chosen ground-state energy $e_{0}$ (this quantity is unknown) for various scenarios. Next, we use the determined in such a way a range for $e_{0}$ and calculate $c(T)$ using $[10,10](e)$ approximant to obtain the curves and shaded areas in the lower panel of Fig. 4. For example, for $\alpha=2$, the values of $e_{0}$ determining the blue curve and shaded area in the lower panel of Fig. 4 are as follows: $e_{0}=-0.439\,2$, $e_{0}=-0.438\,6$, and $e_{0}=-0.437\,9$, cf. the lower panel of Fig. 7.

### v3

@@ 4 Discussion and summary
In this study, we presented the specific heat probes of several quantum three-dimensional Heisenberg magnets. Our main findings are reported in figures 2, 4, and 5. While ferromagnets and bipartite-lattice antiferromagnets can be straightforwardly investigated by applying quantum Monte Carlo simulations, frustrated-lattice antiferromagnets require more sophisticate techniques: We use high-temperature expansion series complemented by the entropy-method interpolation. Our findings illustrate the effects of geometry on the ordering temperature (ferromagnets and bipartite-lattice antiferromagnets). For frustrated-lattice antiferromagnets, the obtained specific heat, which interpolates between the assumed low-energy properties and rigorously known high-temperature ones, provides some information about a possible energy spectrum of the quantum spin system at hand. Although our study is pure theoretical, it may be useful from the solid-state perspective, if the corresponding compounds are available.
@@ Acknowledgements
Appendix A Pyrochlore-lattice antiferromagnet
For the pyrochlore-lattice antiferromagnet, the coefficients $d_{2},\ldots,d_{17}$ are reported in reference [16] (the first thirteen coefficients can be also found in reference [15]). Simple Padé approximants to high-temperature expansion series (2.1) are shown in figure A.1. As can be seen from this figure, they can provide reliable results above $T\approx 0.7$.
To apply the entropy-method interpolation, we use the ground-state energy $e_{0}$ found in references [23, 24] and make an assumption about the decay of the specific heat as $T\to 0$. We try a power-law decay with several exponents $\alpha=1,2,3$, as well as an exponential decay. However, none of the made assumptions yields an excellent temperature profile $c(T)$ with respect to large number $n_{\rm cP}$ of almost coinciding $c(T)$ which follow from $n_{\rm P}=u+d-3=14$ different $[u,d](e)$ approximants. On the other hand, we may try to determine $e_{0}$ as in reference [21]. However, we get $n_{\rm cP}=7$ for $e_{0}=-0.423\,4,\ldots,-0.420\,2$ if $\alpha=1$, $n_{\rm cP}=6$ for $e_{0}=-0.461\,3,\ldots,-0.42$ if $\alpha=2$, or $n_{\rm cP}=7$ for $e_{0}=-0.465\,2,\ldots,-0.457\,8$ if $\alpha=3$; these values are quite far from $e_{0}=-0.49$ [23] and $e_{0}=-0.477$ [24]. Apparently, the calculation of $c(T)$ for the pyrochlore-lattice $S=1/2$ Heisenberg antiferromagnet is far from being finally settled.
Figure A.1: (Colour online) Padé approximants to high-temperature expansion series for the $S=1/2$ Heisenberg antiferromagnet on pyrochlore lattice.
Appendix B Hyperkagome-lattice antiferromagnet
We begin with the coefficients $d_{2},\ldots,d_{16}$ in equation (2.1), which are as follows:
$d_{2}$ $=$ $\frac{3}{8},\quad d_{3}=0,\quad d_{4}=-\frac{51}{128},\quad d_{5}=0,\quad d_{6}=\frac{349}{1\,024},$
$d_{7}$ $=$ $-\frac{7}{1\,280},\quad d_{8}=-\frac{131\,449}{491\,520},$
$d_{9}$ $=$ $\frac{349}{35\,840},\quad d_{10}=\frac{11\,103\,797}{55\,050\,240},$
@@ Appendix B Hyperkagome-lattice antiferromagnet
These coefficients can be found, in principle, in reference [17], where they are presented, although as decimal fractions. In equation (B) they are given as common fractions which may be sometime more convenient.
Figure B.1: (Colour online) Entropy-method calculations for the $S=1/2$ Heisenberg antiferromagnet on hyperkagome lattice. Top: Padé approximants of high-temperature expansion series. Bottom: ground-state energy determination within the entropy method.
The entropy method calculations are described in detail in reference [22], but now we use as an input 20 coefficients in equation (2.1) (instead of 16 in reference [22]). From the upper panel of figure B.1 one concludes that Padé approximants of high-temperature expansion series can give $c(T)$ for $T$ exceeding $\approx 0.3$. In the lower panel of figure B.1 we closely follow reference [22] and present the number of almost coinciding $c(T)$ denoted as $n_{\rm cP}$ among all considered $n_{\rm P}=u+d-3$ Padé approximants $[u,d](e)$ as a function of the chosen ground-state energy $e_{0}$ (this quantity is unknown) for various scenarios. Next, we use the range determined this way for $e_{0}$ and calculate $c(T)$ using $[10,10](e)$ approximant to obtain the curves and the shaded areas in the lower panel of figure 5. For example, for $\alpha=2$, the values of $e_{0}$ determining the blue curve and shaded area in the lower panel of figure 5 are as follows: $e_{0}=-0.4392$, $e_{0}=-0.4386$, and $e_{0}=-0.4379$, cf. the lower panel of figure B.1.